A 4.30 kg block of ice at 0.00 ∘C falls into the ocean and melts. The average temperature of the ocean is 3.55 ∘C, including all the deep water.

Part A:

By how much does the melting of this ice change the entropy of the world?

Part B:

Does it make it larger or smaller?
Answer #1

Part A)

m = mass of ice block = 4.30 kg

L = latent heat of melting = 334000 J/kg

Q = mL

Q = (4.30) (334000)

Q = 1436200 J

Ti = temperature of ice = 273 K

To = temperature of ocean = 3.55 + 273 = 276.55 K

change in entropy is given as

S = Q/Ti - Q/To

S = 1436200/273 - 1436200/276.55

S = 67.5

B)

it makes it larger